OpenStudy (anonymous):

Solve system of equation by addition or substitution method. x=y+2 x+y=8

OpenStudy (usukidoll):

sub the x = y+2 into the second equation

OpenStudy (usukidoll):

y+2+y = 8 2y+2=8 -2 -2 2y = 6 y =3

OpenStudy (usukidoll):

x = 3+ 2 = 5

OpenStudy (anonymous):

Thanks

OpenStudy (usukidoll):

medal?